OpenStudy (anonymous):

Evaluate each expression. 1. 3log34 + log5 25

OpenStudy (btaylor):

Is this \(3 \log_3 4 + \log_5 25\)?

OpenStudy (anonymous):

Yes

OpenStudy (btaylor):

OK...We want to look at a couple properties of logs: \(\log_a b=y \rightarrow a^y = b\) AND \(n \log_a b = log_a \left( b^n \right) \). You can use each of these in your problem.

OpenStudy (mathmale):

@deezyus: Please share with the rest of us what you already know about possible approaches to solving this problem. It's important that you be involved. One suggestion: please double-check that you have copied this problem down correctly. I note that your expression has two logarithms, and that these logs have different bases (3 and 5). That makes more work for us: we need to use the change of base formula. One last thing: some simple examples. There are 3 formulas commonly used with the product, quotient and power functions using logs. One is \[a*\log b=\log b^a\], which btaylor has already introduced. Thus, 2 log 3=log 3^2. Familiar with that?

OpenStudy (anonymous):

I need help solving it

OpenStudy (mathmale):

Another simple example: log a + log b = log a*b

OpenStudy (mathmale):

At this point we need for you to explain what kind of help you need. btaylor and I have given you formulas; I've given you two relatively simple examples. How may I/we help you further?

OpenStudy (anonymous):

I DONT KNOW HOW TO DO IT AT ALL

OpenStudy (mathmale):

Do you have a textbook that shows examples, or at least some online reading materials with examples? Have you been to class where your teacher has discussed solving problems such as these?

OpenStudy (anonymous):

Its a online class no teacher or book

OpenStudy (mathmale):

No examples available in your online reading materials? Have you looked for and read such materials?

OpenStudy (mathmale):

Further examples: log 2 + log 3 = log ( what? )

OpenStudy (mathmale):

Rule: log a - log b = log (a/b) (division rule). Example: log 6 - log 3 = log ( what? )

OpenStudy (anonymous):

I have no accessible materials at all

OpenStudy (mathmale):

Not even on the internet?

OpenStudy (mathmale):

I got the following from http://www.mathwarehouse.com/logarithm/rules-and-formula.php

OpenStudy (mathmale):

Now you have something for reference. These rules are the same as those I've typed out for you. Each rule comes with examples. What do you need to know now, to solve the problem given you?

OpenStudy (anonymous):

Your no help :/